OpenStudy (anonymous):

using the general formula y=mx that describes direct proportionality, find the value of m if y is directlt proportional to x and y=20 when x=1-

OpenStudy (anonymous):

using the general formula y=mx that describes direct proportionality, find the value of m if y is directlt proportional to x and y=20 when x=10****

OpenStudy (tamtoan):

x = 10 ? if yes , you have : y = mx ==> m = y/x = 20/10 = 2 ==> m = 2

OpenStudy (jwheele1):

m is not normally used. m is usually k in classes

OpenStudy (jwheele1):

perhaps this will help? http://www.maths4scotland.co.uk/GHS%20Exam%20Revision/GHS%20Credit/Proportion.swf

OpenStudy (jwheele1):

y = mx 20 = 10m 2=m